OpenStudy (anonymous):

Pablo was graphing a function and noticed that at certain points, the graph reaches invisible lines the graph will never cross. Explain to Pablo what the two types of invisible lines are and how to predict them. You may create your own example to aid in your reasoning. Use complete sentences

OpenStudy (anonymous):

I need help i dont get what its asking

OpenStudy (anonymous):

is anyone good at math cause im not :(

OpenStudy (zzr0ck3r):

look at the graph, of \(f(x)=\frac{1}{x}+1\) notice that as x gets large we get really close to y=1 but never actually cross it.

OpenStudy (anonymous):

it doesnt show me a graph that why im confused :/

OpenStudy (zzr0ck3r):

there is the graph

OpenStudy (anonymous):

oh ok so there 2 invisible line which are?

OpenStudy (zzr0ck3r):

the "invisible" line is the one in purple. notice even if we remove that line, then the black graph gets very close it but never crosses it.

OpenStudy (anonymous):

oh ok how do u predict the invisble line?

OpenStudy (zzr0ck3r):

plug in numbers but this is how you can reason it 1+1/x notice as x gets very big \(\frac{1}{x}\) gets close to 0, but never actually hits 0 so \(1+\frac{1}{x}\) gets close to 1, but never crosses it.

OpenStudy (anonymous):

ok thanks and im sorry math is my worse subject :(

OpenStudy (zzr0ck3r):

np:)

OpenStudy (anonymous):

could i ask you antheir question